### v1

@@ 4.2 Bubble detachment
The removal of growing bubbles is crucial for enhancing the efficiency of the electrolysis process (Angulo et al., 2020). The bubble detachment radius is approximately equal to the Fritz radius (Stephan and Vogt, 1979)
$R_{det}=0.6\theta\sqrt{\frac{\sigma}{(\rho_{c}-\rho_{d})g}},$ (37)
where $\theta$ the contact angle in radians. The general contact angle for the bubble at the electrode would be smaller than $\pi/2$. This analytical solution is derived by balancing the buoyancy force for a perfect sphere-shaped bubble $F_{b}=(4/3)\pi R^{3}(\rho_{c}-\rho_{d})g$ with the capillary force for a bubble attached to the electrode. The contact angle controls the magnitude of surface tension effects and varies in connection with contact line dynamics, which are still, in many circumstances, an object of investigation. Molecular dynamics simulations with either Lennard-Jones intermolecular potentials or more realistic water molecule models give very different results, see Lācis et al. (2022). Therefore, in the present study, the theoretical value is adopted as a reference to estimate the order of magnitude rather than for precise comparison.
@@
To begin with, a simulation of single-bubble detachment is shown in fig. 16. The numerical configuration refers to the No.1 and No.4 in Table 1. As observed, once the mass transfer is considered (at $t=20\mathrm{ms}$), the bubble starts to grow. Initially, the bubble remains nearly spherical due to its small size, which results in a negligible buoyancy force. However, as the bubble radius increases, deformation becomes evident, particularly after $t=52.1\mathrm{ms}$. Eventually, buoyancy overcomes surface tension, leading to detachment from the electrode surface.
Figure 16: Frames show the detachment evolution($\theta=90^{\circ}$).
As previously mentioned, eq. 37 establishes a relationship between the detachment radius and the contact angle. To further investigate this dependency, another case with a contact angle $\theta=35\degree$ was also simulated, see fig. 17. Compared to the case $\theta=90\degree$, the detachment time is much shorter, and the detachment radius is much smaller. fig. 18 presents a good agreement between the numerical simulation and the Fritz radius.
Figure 17: Frames show the detachment evolution($\theta=35^{\circ}$).
@@
Figure 18: Bubble detachment radius versus contact angle for a single bubble
@@
Additionally, it has been confirmed that mutual interactions between bubbles influence the detachment process. Bashkatov et al. (2024) reported that bubble coalescence in the presence of dual bubbles can lead to significantly earlier detachment and a smaller detachment radius compared to cases where only buoyancy effects are considered. To validate this finding, we perform a series of simulations involving bubble detachments for multiple nucleation sites. The results of 3D simulations with four nucleation sites are presented in fig. 19. Furthermore, varying the number of nucleation sites demonstrates that bubble coalescence indeed accelerates detachment, as shown in fig. 20. As a result, the bubble detachment radius is also smaller, see fig. 21.
A plausible mechanism behind this acceleration could be the following. The coalescence-induced shape perturbations generate an initial momentum favoring premature detachment from the electrode surface. It is observed in both our simulation and Bashkatov’s experiment that the coalescence of bubbles results in an initial jump-off of the merged bubble. This can be attributed to the released surface energy during coalescence (Zhang et al., 2024). The released energy is partly dissipated by the bubble oscillations, working against viscous drag. When in the proximity to the surface, the remaining energy is converted to kinetic energy, driving the resultant (merged) bubble to jump off the electrode (Lv et al., 2021). This kinematic energy can cause bubble departure at smaller radii than in the buoyancy-driven scenario, as observed in our simulation results. However, the capillary wave due to energy dissipation is not observed in our simulation because of the mesh resolution constraint. Further numerical investigations are needed to distinguish the respective roles of buoyancy and interfacial effects in this process.
Figure 19: snapshots of the detachment evolution for 4 nucleation sites (Configuration No.9).
@@
Figure 20: Bubble detachment times for different nucleation sites and contact angles.
@@
Figure 21: Bubble detachment radius for different nucleation sites and contact angles.
@@ 5 Conclusion
Bubble growth dynamics The bubble contact angle significantly influences growth behavior by dictating the geometric position of the bubble within the evolving hydrogen concentration profile. This directly modulates local mass transport rates. For multi-bubble systems, mutual suppression of mass transport occurs due to simultaneous hydrogen consumption by neighboring bubbles, as quantified by reductions in Sherwood number (fig. 15). While simulations overpredict growth rates compared to experimental observations (fig. 13), they adhere closely to the scaling law of Scriven’s solution ($R\propto t^{1/2}$).
Bubble detachment and coalescence behavior Detachment radii align with Fritz’s formula, demonstrating a decrease for smaller contact angles. Multi-bubble systems exhibit earlier detachment with reduced radii compared to isolated bubbles. t is worth noting that coalescence reduces the detachment time, thus potentially improving the mass transfer efficiency of the electrochemical system.
Prospective The present work focuses on single growth-detachment cycles due to computational constraints. Tracking deformable interfaces remains resource-intensive, limiting simulations to short time spans ($t\sim 0.2$s). While comparable to prior studies using rigid bubble models (Sepahi et al., 2022), broader parametric investigations (e.g., current density variations, contact angle distributions) are needed to strengthen statistical conclusions. Future efforts are needed to develop a new scheme with subgrid model, which will help to enable long-term simulations.
Additionally, it should be noted that the experimental data used for comparison are not fully aligned with the study’s objectives. We will design and implement a more sophisticated experimental setup that will provide observational proof of our work (e.g., continuous single bubble detachment), thus increasing the reliability and impact of the 3D numerical study.

### v4

@@ 4.3 Bubble detachment
The removal of attached bubbles is crucial for enhancing the efficiency of the electrolysis process (Angulo2020). In the absence of contact angle hysteresis, the bubble tends to spread on the electrode surface with constant speed. The force balance determines the relation between the volume-equivalent detachment radius (Fritz radius) and the value of the contact angle $\theta$ giving(Stephan1979),
$R_{det}=0.6\,\theta\sqrt{\frac{\sigma}{(\rho_{c}-\rho_{d})g}}.$ (37)
This analytical solution is derived by balancing the buoyancy force for a perfect sphere-shaped bubble $F_{b}=(4/3)\pi R^{3}(\rho_{c}-\rho_{d})g$ with the capillary force for a bubble attached to the electrode.
@@
To begin with, a simulation of single-bubble detachment is performed. Initially, the bubble remains nearly spherical due to its small size, which results in a negligible buoyancy force. However, as the bubble radius increases, deformation becomes evident. Eventually, buoyancy overcomes surface tension, leading to the detachment from the electrode surface.
@@
Figure 14: Bubble detachment radius versus contact angle for a single bubble(with the current density of $I=\mathrm{1000A/m^{2}}$).
Figure 15: Frames show the detachment evolution for a single bubble at different contact angles(with the current density of $I=\mathrm{1000A/m^{2}}$). The corresponding detachment times ($t_{d}$) are $104.2\mathrm{ms}$ for $\theta=35^{\circ}$, and $243.2\mathrm{ms}$ for $\theta=90^{\circ}$. The hydrogen contour on the intersecting plane (yz) passing through the middle of the bubble is projected to the back plane. The bubble interface is colored according to the local mass-transfer rate.
As previously mentioned, Eq.˜37 establishes a linear relationship between $R_{d}$ and $\theta$. To further investigate this dependency, a series of cases with varying contact angle from $\theta=35\degree$ to $\theta=90\degree$ are simulated, see Fig.˜14. As expected, the numerical results also exhibit a linear relationship. The Fritz radius is computed using the same reduced surface tension. However, the simulation predicts larger detachment radii compared to the theoretical value. The pinning effect of contact angle hysteresis can explain this deviation. For detachment to occur, the contact line must recede (shrink inwards). However, contact angle hysteresis pins the contact line. The bubble cannot simply shrink at its base; instead, its neck must stretch and thin until the surface tension force can no longer hold it. Since the theoretical Fritz radius is derived from a simplified force balance that neglects contact angle hysteresis, the theoretical value is expected to be smaller than the detachment radius predicted by the simulation.
To demonstrate the effect of contact angle on detachment radius, we take the most contrasting cases, $\theta=90^{\circ}$ (hydrophobic) and $\theta=35^{\circ}$ (hydrophilic), as an example. As shown in Fig.˜15, the bubble with the smaller contact angle ($\theta=35^{\circ}$) detaches significantly faster and at a much smaller radius.
@@
Additionally, it has been shown that mutual interactions between bubbles influence the detachment process. Bashkatov2024 observed that bubble coalescence in the presence of dual bubbles can lead to significantly earlier detachment and a smaller detachment radius compared to cases where only buoyancy effects are considered. As illustrated in Fig.˜16, the presence of multiple bubbles indeed accelerates the detachment. Consequently, the detachment radius is reduced, as demonstrated in Fig.˜17.
For a direct visual comparison, representative frames from simulations with multiple bubbles at contact angles of $\theta=90^{\circ}$ and $\theta=35^{\circ}$ are provided in Figs.˜18 and 19.
@@
Figure 16: Bubble detachment times for different nucleation sites and contact angles (with the current density of $I=\mathrm{1000A/m^{2}}$).
@@
Figure 17: Bubble detachment radius for different nucleation sites and contact angles (with the current density of $I=\mathrm{1000A/m^{2}}$).
Figure 18: Frames show the detachment evolution for multi-bubbles at the contact angle of $\theta=90\degree$ (with the current density of $I=\mathrm{1000A/m^{2}}$). The corresponding detachment times ($t_{d}$) are $243.2\mathrm{ms}$ for 1 bubble, $241.4\mathrm{ms}$ for the 2-bubbles case, and $235.5\mathrm{ms}$ for the 4-bubbles case. The hydrogen contour on the intersecting plane (yz) passing through the middle of the bubble is projected to the back plane. The bubble interface is colored according to the local mass-transfer rate.
Figure 19: Frames show the detachment evolution for multi-bubbles at the contact angle of $\theta=35\degree$ (with the current density of $I=\mathrm{1000A/m^{2}}$). The corresponding detachment times ($t_{d}$) are $104.2\mathrm{ms}$ for 1 bubble, $94.6\mathrm{ms}$ for the 2-bubbles case, and $67.6\mathrm{ms}$ for the 4-bubbles case. The hydrogen contour on the intersecting plane (yz) passing through the middle of the bubble is projected to the back plane. The bubble interface is colored according to the local mass-transfer rate.
A plausible mechanism behind this acceleration is as follows. The coalescence-induced shape perturbations generate an initial momentum favoring premature detachment from the electrode surface. It is observed in Bashkatov’s experiment that the coalescence of bubbles results in an initial jump-off of the merged bubble. This can be attributed to the released surface energy during coalescence (Zhang2024). The released energy is partly dissipated by the bubble oscillations, working against viscous drag. When in the proximity to the surface, the remaining energy is converted to kinetic energy, driving the resultant (merged) bubble to jump off the electrode (lv2021). This kinetic energy can cause bubble departure at smaller radii than in the purely buoyancy-driven scenario, as observed in our simulation results. However, the capillary waves due to energy dissipation are not observed in our simulations because of the mesh resolution constraint. Further numerical investigations are needed to distinguish the respective roles of buoyancy and interfacial effects in this process.
@@ 5 Conclusion
Bubble growth dynamics
The bubble contact angle significantly influences growth behavior by dictating the geometric position of the bubble within the evolving hydrogen concentration profile. This geometric position directly modulates the local mass-transfer rates.
For multi-bubble systems, mutual suppression of mass transport occurs due to the simultaneous hydrogen consumption by neighboring bubbles, a phenomenon quantified by reductions in the Sherwood number ($\mathrm{Sh}_{b}$), as shown in Fig.˜11. Specifically, increasing the number of nucleation sites leads to a smaller individual bubble Sherwood number. While some numerical overprediction and underprediction of growth rates exist compared to experimental observations (Fig.˜12), the simulations adhere closely to the scaling law of Scriven’s solution ($R\propto t^{1/2}$).
Bubble detachment and coalescence behavior
Detachment radii align with Fritz’s formula, demonstrating a linear relation between the volume-equivalent radii and the contact angles. Multi-bubble cases exhibit earlier detachment with reduced radii compared to isolated bubbles. It is worth noting that coalescence reduces the detachment time, thus potentially improving the mass transfer efficiency of the electrochemical system.
Outlook
The present work focuses on single growth-detachment cycles due to computational constraints. Tracking deformable interfaces remains resource-intensive, limiting simulations to short time spans ($t\sim 0.2\,\mathrm{s}$). While comparable to prior studies using rigid bubble models (sepahi2022a), broader parametric investigations (e.g., current density variations, contact angle distributions) are necessary to strengthen statistical conclusions. For instance, a statistically steady Sherwood number at the electrode $\mathrm{Sh_{e}}$ has not been investigated in this work. Future efforts are required to develop a new scheme with a larger stable time step, which will help to enable long-term simulations.